OpenStudy (anonymous):

Solve the following system of equations for x: x - 7y = - 14 6x - 8y = 18

OpenStudy (anonymous):

A. None of these. B. x = - 3 C. x = 3 D. x = 7 E. x = .5

OpenStudy (anonymous):

x - 7y = - 14 ---(1) 6x - 8y = 18 --- (2) from (1), x=7y-14--(3) sub it into (2), 6(7y-14) -8y =18 42y-84 -8y =18 34y=102 y=3 put it into (3) x=7(3)-14 = 7

OpenStudy (anonymous):

Hi @vince33 - it looks like a lot of your questions seem to be on the same type of question (system of equations). My suggestion would be to ask one of them, and then see if you can use what you've learned to answer the others.